by examining the mean amount of bacteria present per plate of culture when a fixed amount of the antiblotic is applied. Previous exper with the antiblotic on this type of bacteria Indicates that the standard deviation of the amount of bacteria present is approximately 19 cm2 Use this to determi e the nur ber of obervations (cultures that must be developed and then tested) to estimate the mean amount of bacteria present, using a 99% confidence interval with a half width of 4 cm2
Answer #1

Solution:

Given in the question

Standard deviation = 19

Alpha =0.01

Alpha/2 = 0.005 so Zalpha/2 = 2.575

Margin of error = 4

So margin of error can be calculated as

Margin of error = Zalpha/2*SD/sqrt(n)

4 = 2.575*19/sqrt(n)

N = (2.575*19/4)^2 = (12.23125)^2 = 150